etween June and July rease in the cost per equivalent unit for direct matenials and PR 17-3A Equivalent units and related costs; cost entries White Diamond Flour Com of production report. Obi, 2,3,4 2. Transferred to packagi $40,1s3 materials pass throusgh Dept., W Dept. ith wheat grain being introduced pany manufactures flour by a series of three processes, beginning ced in the Milling Department. From the Milling Department, the packaged refined flour was as follows on July 1 s pass through the Sifting and Packaging departments, emerging as Milling Depa The balance in the account Work in Process-Sifting Department Work in Process-Sifting Department (900 units, /scompleted) Direct materials (900 × $2.05) Conversion (900 × ⅓ × S040) EXCEL TEMPLATE $1,845 216 S2,061 The following costs were charged to Work in Process-si fting Department during Julsy Direct materials transferred from Milling Department $33,755 15,700 units at $2.15 a unit Direct labor Factory overhead 4,420 2,708 s-Sifting Department on During July, 15,500 units of flour were completed. Work in Proces July 31 was 1,100 units, ts completed. Instructions 1. Prepare a cost of production report for the Sifting Department for July 2. Journalize the entries for costs transferred from Milling to Sifting and the costs transferr uly for from Sifting to Packaging 3. Determine the increase or decrease in the cost per equivalent unit from June to direct materials and conversion costs. 4. Discuss the uses of the cost of production report and the results of part (G3).
Answer #1

1)

WHITE DIAMOND FLOUR COMPANY
Cost of Production Report
Units Complete Units Equivalent Units
Direct Material Conversion
Units charged to production:
Opening Inventory in Process                    900.00
Units from Milling Department              15,700.00
Total Units              16,600.00
Units to be assigned Cost
Opening Inventory in Process                    900.00                     -              360.00
Units completed in July              14,600.00      14,600.00      14,600.00
Transfer to Packaging Dept              15,500.00      14,600.00      14,960.00
Ending Inventory in Process (4/5)                1,100.00        1,100.00            880.00
Total Units to be assigned Cost              16,600.00      15,700.00      15,840.00
Costs Costs
Direct Material Conversion Total
Total Cost $          33,755.00 $    7,128.00
Equivalent units              15,700.00      15,840.00
Cost per Equivalent Unit $                    2.15 $            0.45
Cost allocated to Production
For Opening Inventory in Process $    2,061.00
Cost in July (33,755 + 4,420 + 2,708) $ 40,883.00
Total Cost $ 42,944.00
Cost allocated to completed and partially completed units:
For Opening Inventory in Process $    2,061.00
To complete inventory in process $        162.00 $        162.00
Total $    2,223.00
Started and completed in July 31390 6570 37960
Transfer to Packaging Dept $ 40,183.00
Ending Inventory In Process 2365 396 $    2,761.00
Total Cost Allocated to Sifting Department $ 42,944.00

2)

Debit Credit
Work in Process—Sifting Department $ 33,755.00
Work in Process—Milling Department $ 33,755.00
Work in Process—Packaging Department   $ 40,183.00
Work in Process—Sifting Department $ 40,183.00

3)Direct Material Increase in Cost = 2.15 - 2.05 = 0.10

Conversion = 0.45 - 0.40 = 0.05

4)The cost of Production report helps in assigning costs between work in Process and also the finished goods . It also helps in controlling cost Department wise. Any changes in Cost can be analysed as done in Part 3 and proper action can be taken.